# From Pixels to Words – Towards Native Vision-Language Primitives at Scale

### v1

@@ 4.1 Training Settings
Our NEO models are built on Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the LLMs. The pre-Buffer employs $L_{1}=$ 12 primitive layers for NEO-2.2B and $L_{1}=$ 6 for NEO-9B. We extend only the QK head dimension in raw transformer layers, introducing roughly 10% extra parameters over the original design. The base RoPE frequencies $\beta_{T}$, $\beta_{H}$, and $\beta_{W}$ are set to $\times 10^{6}, 1×1$$\times 10^{4}$, and $1\times 10^{4}$, respectively. NEO is trained on sixteen 8-GPU (80G) nodes using the AdamW optimizer Loshchilov & Hutter (2019). The maximum learning rates for pre-training, mid-training, and SFT are $8\times 10^{-4}$, $4\times 10^{-5}$, and $5\times 10^{-5}$, with a warm-up ratio of 0.01 and a cosine decay scheduler across all stages.
@@ 4.2 Main Results
We conduct standard evaluations with VLMEvalKit Duan et al. (2024) on diverse benchmarks, covering chart, diagram, and document understanding tasks, e.g., AI2D Kembhavi et al. (2016), DocVQA Clark & Gardner (2018), ChartQA Masry et al. (2022), InfoVQA Mathew et al. (2022), TextVQA Singh et al. (2019), and OCRBench Liu et al. (2023e); visual perception and challenging reasoning tasks, e.g., MMMU Yue et al. (2024), MMBench-EN (MMB) Liu et al. (2024b), MMVet Yu et al. (2024), MMStar Chen et al. (2024c), SEEDBench-IMG (SEED-I) Li et al. (2023a); visual hallucination tasks, e.g., POPE Li et al. (2023b) and HallusionBench (HallB) Guan et al. (2024).
@@
Following InternVL3 Zhu et al. (2025), we construct the Encoder-Based by combining Qwen3 Yang et al. (2025) and InternViT-300M Zhu et al. (2025). In the mid-training stage, we first train the projector on 10M samples, and further unfreeze the vision encoder utilizing another 30M samples.
Comparison with Modular VLMs. As demonstrated in Table 1 and Table 2, NEO achieves highly competitive performance at both the 2B and 8B scales, despite using relatively limited pre-training and supervised fine-tuning data and without reinforcement learning. Remarkably, NEO approaches the performance of top-tier modular VLMs, e.g., Qwen2-VL Wang et al. (2024a), InternVL2.5 Chen et al. (2024e), Qwen2.5-VL Bai et al. (2025), and InternVL3 Zhu et al. (2025) across multiple benchmarks, rivaling architectures trained on billions of additional samples. These results highlight the effectiveness of our end-to-end training strategy and unified model design. By combining native attention mechanisms with Native-RoPE, NEO enhances interactions between visual and linguistic features, enabling it to match more complex modular systems despite its simpler architecture.
Comparison with Native VLMs. From Table 1 and Table 2, NEO delivers substantial gains on visual-centric benchmarks over the best competitors, e.g., Mono-InterVL Luo et al. (2024; 2025), HoVLE Tao et al. (2025), OnCAT Li et al. (2025a), EVE Diao et al. (2024; 2025), Emu3 Wang et al. (2024b), BREEN Li et al. (2025b), VoRA Wang et al. (2025a), and SAIL Lei et al. (2025). By seamlessly integrating post-LLM components with the pre-Buffer for large-scale visual learning, NEO aligns visual inputs with textual features from scratch and supports complex visual reasoning, even without visual encoder supervision Diao et al. (2024); Tao et al. (2025); Li et al. (2025a); Wang et al. (2025a); Li et al. (2025b), highlighting the strengths of its native primitive designs and training strategies. These design choices allow NEO to surpass many native VLMs using fewer training resources, demonstrating the advantages of our primitives with efficient data-scaling capability.
@@ 4.3 Ablation Studies
Hyperparameters of the Pre-Buffer Layer. Figure 5 illustrates the relationship between the number of pre-Buffer layers and the model’s average accuracy, using Qwen3-1.7B as the post-LLM. Performance improves consistently as the layer count increases, but gains begin to saturate beyond eight layers. To maximize accuracy while maintaining the same capacity as publicly available vision encoders Chen et al. (2024f); Radford et al. (2021); Zhai et al. (2023), we select 12 layers for NEO-2.2B. Notably, we choose 6 layers for NEO-9B, mainly due to the good trade-off between performance and efficiency.
@@
Figure 6: Comparison with pre-Buffer and vision encoders. All models are initialized post-LLM using Qwen3-1.7B.
Figure 7: Evaluation results across three progressive training procedures.
Configurations of Native Primitives. Table 3 compares various attention and RoPE designs. The pre-Buffer depth is 4, and the post-LLM is initialized with Qwen3-1.7B. All models share the same new QK head dimensions and normalization. (1) Attention mode. Comparing models A/B and G/H reveals consistent gains of mixed attention over causal one, reflecting its stronger capacity to model comprehensive dependencies and cross-modal alignment. (2) RoPE mode. Native-RoPE outperforms 1D-RoPE Zhu et al. (2025), IL-RoPE Liao et al. (2025), M-RoPE Bai et al. (2025), MM-RoPE Yuan et al. (2025), and Video-RoPE Wei et al. (2025), with at least a 0.8% gain. This validates the importance of disentangling height, width, and temporal components in RoPE to enhance spatial–temporal representations and fine-grained interactions. By contrast, setting the base RoPE frequency to 1M for height and width severely impairs the ability to perceive local semantics.
Comparison between Pre-Buffer and Vision Encoders. In Figure 7, PB 1–3 denotes the Pre-Buffer after stage 1–3. For all models except NEO, the post-LLMs are initialized via Qwen3-1.7B for our pre-Buffer, InternViT-300M Chen et al. (2024e), CLIP-vit-large-patch14 Radford et al. (2021), and SigLIP-so400m-patch14-384 Zhai et al. (2023). After two-stage re-training, PB3 shows only an average gap of 2.5 / 2.4 / 1.7 / 3.7% over NEO / InternViT / CLIP / SigLIP using billion-scale training data. This substantially reduces the training costs of building native VLMs for subsequent research.
@@ A.1 Supervised Fine-tuning Datasets
Captioning TextCaps (en) Sidorov et al. (2020), ShareGPT4V (en&zh) Chen et al. (2024b)
VQAv2 (en) Goyal et al. (2017), GQA (en) Hudson & Manning (2019), OKVQA (en) Marino et al. (2019),
General QA VSR (en) Liu et al. (2023a), VisualDialog (en) Das et al. (2017)
Science AI2D (en) Kembhavi et al. (2016), ScienceQA (en) Lu et al. (2022a), TQA (en) Kembhavi et al. (2017)
ChartQA (en) Masry et al. (2022), MMC-Inst (en) Liu et al. (2023c), DVQA (en) Kafle et al. (2018),
Chart PlotQA (en) Methani et al. (2020), LRV-Instruction (en) Liu et al. (2023b)
GeoQA+ (en) Cao & Xiao (2022), TabMWP (en) Lu et al. (2022b), MathQA (en) Yu et al. (2023),
Mathematics CLEVR-Math/Super (en) Lindström & Abraham (2022); Li et al. (2023c), Geometry3K (en) Lu et al. (2021)
KVQA (en) Shah et al. (2019), A-OKVQA (en) Schwenk et al. (2022), ViQuAE (en) Lerner et al. (2022),
Knowledge Wikipedia (en&zh) He et al. (2023)
OCRVQA (en) Mishra et al. (2019), InfoVQA (en) Mathew et al. (2022), TextVQA (en) Singh et al. (2019),
ArT (en&zh) Chng et al. (2019), COCO-Text (en) Veit et al. (2016), CTW (zh) Yuan et al. (2019),
LSVT (zh) Sun et al. (2019), RCTW-17 (zh) Shi et al. (2017), ReCTs (zh) Liu et al. (2019),
OCR SynthDoG (en&zh) Kim et al. (2022), ST-VQA (en) Biten et al. (2019)
Document DocVQA (en) Clark & Gardner (2018), Common Crawl PDF (en&zh)
Grounding RefCOCO/+/g (en) Yu et al. (2016); Mao et al. (2016), Visual Genome (en) Krishna et al. (2017)
LLaVA-150K (en&zh) Liu et al. (2023d), LVIS-Instruct4V (en) Wang et al. (2023),
ALLaVA (en&zh) Chen et al. (2024a), Laion-GPT4V (en) LAION (2023),
Conversation TextOCR-GPT4V (en) Jimmycarter (2023), SVIT (en&zh) Zhao et al. (2023)
OpenHermes2.5 (en) Teknium (2023), Alpaca-GPT4 (en) Taori et al. (2023), COIG-CQIA (zh) Bai et al. (2024),
Text-only ShareGPT (en&zh) Zheng et al. (2023)
@@ A.3 Limitation and Discussion
Constrained by current text corpus and computational resources, we are unable to train a fully native model entirely from scratch without initialization from an existing LLM. This limitation also hinders our ability to mitigate potential biases arising from the dominance of the language modality. Despite these challenges, our NEO extends beyond providing a reusable pre-buffer that lowers the cost of adapting advanced LLMs—with updated weights and stronger capabilities—into VLMs under limited budgets. More importantly, NEO reveals the potential performance ceiling of native VLM architectures and provides valuable insights for future research on de novo multimodal training.

### v2

@@ 4.1 Training Settings
Our NEO models are built on Qwen3-1.7B and Qwen3-8B (Yang et al., 2025) as the LLMs. The pre-Buffer employs $L_{1}=$ 12 primitive layers for NEO-2.2B and $L_{1}=$ 6 for NEO-9B. We extend only the QK head dimension in raw transformer layers, introducing roughly 10% extra parameters over the original design. The base RoPE frequencies $\beta_{T}$, $\beta_{H}$, and $\beta_{W}$ are set to $\times 10^{6}, 1×1$$\times 10^{4}$, and $1\times 10^{4}$, respectively. NEO is trained on sixteen 8-GPU (80G) nodes using the AdamW optimizer (Loshchilov and Hutter, 2019). The maximum learning rates for pre-training, mid-training, and SFT are $8\times 10^{-4}$, $4\times 10^{-5}$, and $5\times 10^{-5}$, with a warm-up ratio of 0.01 and a cosine decay scheduler across all stages.
@@ 4.2 Main Results
We conduct standard evaluations with VLMEvalKit (Duan et al., 2024) on diverse benchmarks, covering chart, diagram, and document understanding tasks, e.g., AI2D (Kembhavi et al., 2016), DocVQA (Clark and Gardner, 2018), ChartQA (Masry et al., 2022), InfoVQA (Mathew et al., 2022), TextVQA (Singh et al., 2019), and OCRBench (Liu et al., 2023e); visual perception and challenging reasoning tasks, e.g., MMMU (Yue et al., 2024), MMBench-EN (MMB) (Liu et al., 2024b), MMVet (Yu et al., 2024), MMStar (Chen et al., 2024c), SEEDBench-IMG (SEED-I) (Li et al., 2023a); hallucination tasks, e.g., POPE (Li et al., 2023b) and HallusionBench (HallB) (Guan et al., 2024).
@@
Following InternVL3 (Zhu et al., 2025), we construct the Encoder-Based by combining Qwen3 (Yang et al., 2025) and InternViT-300M (Zhu et al., 2025). In the mid-training stage, we first train the projector on 10M samples, and further unfreeze the vision encoder utilizing another 30M samples.
Comparison with Modular VLMs. In Table 1 and Table 2, NEO achieves highly competitive performance against Encoder-Based counterparts at the 2B and 8B scales. Impressively, NEO largely narrows the performance gap with top-tier modular VLMs, e.g., Qwen2-VL (Wang et al., 2024a), InternVL2.5 (Chen et al., 2024e), Qwen2.5-VL (Bai et al., 2025), and InternVL3 (Zhu et al., 2025) across multiple benchmarks, despite using relatively limited training data and without reinforcement learning. These results highlight the effectiveness of an end-to-end training strategy and a unified model design with Native-RoPE and multi-modality interaction patterns. Moreover, the performance gap between Encoder-Based variants and state-of-the-art methods on MMMU, MMVet, TextVQA, and etc, indicates that NEO still suffers from the limitation in training data scale and quality.
Comparison with Native VLMs. From Table 1 and Table 2, NEO delivers substantial gains on visual-centric benchmarks over the best competitors, e.g., Mono-InterVL (Luo et al., 2024; 2025), HoVLE (Tao et al., 2025), OneCAT (Li et al., 2025a), EVE (Diao et al., 2024; 2025), Emu3 (Wang et al., 2024b), BREEN (Li et al., 2025b), VoRA (Wang et al., 2025a), and SAIL (Lei et al., 2025). By seamlessly integrating post-LLM components with the pre-Buffer for large-scale visual learning, NEO aligns visual inputs with textual features from scratch and supports complex visual reasoning, even without visual encoder supervision (Diao et al., 2024; Tao et al., 2025; Li et al., 2025a; Wang et al., 2025a; Li et al., 2025b), highlighting the strengths of its native primitive designs and training strategies. These design choices allow NEO to surpass many native VLMs using fewer training resources, demonstrating the advantages of our primitives with efficient data-scaling capability.
@@ 4.3 Ablation Studies
Hyperparameters of the Pre-Buffer Layer. Figure 5 illustrates the relationship between the number of pre-Buffer layers and the model’s average accuracy, using Qwen3-1.7B as the post-LLM. Performance improves consistently as the layer count increases, but gains begin to saturate beyond eight layers. Inspired by the parameter counts of popular VEs (Chen et al., 2024f; Radford et al., 2021; Zhai et al., 2023) and the scaling properties (Tian et al., 2025) between existing VEs and LLMs, we select 12 and 6 layers for NEO-2.2B and NEO-9B to balance the trade-off between performance and efficiency.
@@
Figure 6: Pre-Buffer vs. VEs.
Figure 7: Neo vs. EVE series.
Figure 8: Three training processes.
Configurations of Native Primitives. Table 3 compares various attention and RoPE designs. The pre-Buffer depth is 4, and the post-LLM is initialized with Qwen3-1.7B. All models share the same new QK head dimensions and normalization. (1) Attention mode. Comparing models A/B and G/H reveals consistent gains of mixed attention over causal one, reflecting its stronger capacity to model comprehensive dependencies and cross-modal alignment. (2) RoPE mode. Native-RoPE outperforms 1D-RoPE (Zhu et al., 2025), IL-RoPE (Liao et al., 2025), M-RoPE (Bai et al., 2025), MM-RoPE (Yuan et al., 2025), and Video-RoPE (Wei et al., 2025), with at least a 0.8% gain. This validates the importance of disentangling height, width, and temporal components in RoPE to enhance spatial–temporal representations and fine-grained interactions. By contrast, setting the base RoPE frequency to 1M for height and width severely impairs the ability to perceive local semantics.
Comparison between Pre-Buffer and Vision Encoders. In Figure 8, PB 1–3 denotes the Pre-Buffer obtained from stage 1–3. For fairness, all post-LLMs are initialized by Qwen3-1.7B (Yang et al., 2025) combined by pre-trained pre-Buffer, CLIP-vit-large-patch14 (Radford et al., 2021), InternViT-300M (Chen et al., 2024e), and SigLIP-so400m-patch14-384 (Zhai et al., 2023). During pre-training, we train the projector for encoder-based methods and the newly added QK parameters for all models. During SFT, we unfreeze the entire backbone. After two-stage re-training, PB3 delivers strong performance, trailing CLIP / InternViT / SigLIP by only 1.7 / 2.4 / 3.7% on average across diverse benchmarks. Notably, despite using only 22M training samples, PB3 is just 2.5% below the full NEO model, substantially reducing the training cost of developing native VLMs for future research.
Comparison between NEO and Native VLM variants. Existing native VLMs in Table 1 and 2 differ dramatically in training pipelines, data volume/quality, and base LLM choice. To isolate model architectural effects, we compare our NEO with representative EVEv1.0 (Dense), EVEv1.5 (Mixture-of-Experts), and EVEv2.0 (Divide-and-Conquer), all built on Qwen3-1.7B. From Figure 8, EVEv1.0, EVEv1.5, and EVEv2.0 achieves 33.4%, 40.2%, and 41.5%, respectively, compared with 44.0% for our NEO. This confirms that the gains stem from our pre-Buffer, native RoPE design, and interaction patterns, rather than from a newer backbone or larger and higher-quality data alone.
@@ A.1 Supervised Fine-tuning Datasets
Captioning TextCaps (en) (Sidorov et al., 2020), ShareGPT4V (en&zh) (Chen et al., 2024b)
VQAv2 (en) (Goyal et al., 2017), GQA (en) (Hudson and Manning, 2019), OKVQA (en) (Marino et al., 2019),
General QA VSR (en) (Liu et al., 2023a), VisualDialog (en) (Das et al., 2017)
Science AI2D (en) (Kembhavi et al., 2016), ScienceQA (en) (Lu et al., 2022a), TQA (en) (Kembhavi et al., 2017)
ChartQA (en) (Masry et al., 2022), MMC-Inst (en) (Liu et al., 2023c), DVQA (en) (Kafle et al., 2018),
Chart PlotQA (en) (Methani et al., 2020), LRV-Instruction (en) (Liu et al., 2023b)
GeoQA+ (en) (Cao and Xiao, 2022), TabMWP (en) (Lu et al., 2022b), MathQA (en) (Yu et al., 2023),
Mathematics CLEVR-Math/Super (en) (Lindström and Abraham, 2022; Li et al., 2023c), Geometry3K (en) (Lu et al., 2021)
KVQA (en) (Shah et al., 2019), A-OKVQA (en) (Schwenk et al., 2022), ViQuAE (en) (Lerner et al., 2022),
Knowledge Wikipedia (en&zh) (He et al., 2023)
OCRVQA (en) (Mishra et al., 2019), InfoVQA (en) (Mathew et al., 2022), TextVQA (en) (Singh et al., 2019),
ArT (en&zh) (Chng et al., 2019), COCO-Text (en) (Veit et al., 2016), CTW (zh) (Yuan et al., 2019),
LSVT (zh) (Sun et al., 2019), RCTW-17 (zh) (Shi et al., 2017), ReCTs (zh) (Liu et al., 2019),
OCR SynthDoG (en&zh) (Kim et al., 2022), ST-VQA (en) (Biten et al., 2019)
Document DocVQA (en) (Clark and Gardner, 2018), Common Crawl PDF (en&zh)
Grounding RefCOCO/+/g (en) (Yu et al., 2016; Mao et al., 2016), Visual Genome (en) (Krishna et al., 2017)
LLaVA-150K (en&zh) (Liu et al., 2023d), LVIS-Instruct4V (en) (Wang et al., 2023),
ALLaVA (en&zh) (Chen et al., 2024a), Laion-GPT4V (en) (LAION, 2023),
Conversation TextOCR-GPT4V (en) (Jimmycarter, 2023), SVIT (en&zh) (Zhao et al., 2023)
OpenHermes2.5 (en) (Teknium, 2023), Alpaca-GPT4 (en) (Taori et al., 2023), COIG-CQIA (zh) (Bai et al., 2024),
Text-only ShareGPT (en&zh) (Zheng et al., 2023)